Week of 151116

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Maarten (ALICE), Fernando (Batch), Andrei (DB)
  • remote: Dario (ATLAS), Giuseppe (CMS), Raja (LHCb), Michael (BNL), Francesco (CNAF), Rolf (IN2P3), Sang Un (KISTI), Pavel (KIT), Onno (NL-T1), Chris (OSG), Jose (PIC), John (RAL), Di Qing (TRIUMF)

Experiments round table:

  • CMS reports (raw view) -
    • No major issues to report
    • Preparing for Heavy Ion run

  • ALICE -
    • Our thanks go to a number of big sites for setting up high-memory queues for heavy ion data reconstruction!
    • KISTI, KIT and NLT1 are ready
    • CERN and CNAF are in progress

  • LHCb reports (raw view) -
    • Data Processing:
      • Data processing of pp data at T0/T1/T2 sites.
      • Monte Carlo mostly at T0/T1/T2/T2D, user analysis at T0/1/2D sites
    • T0
      • NTR
    • T1
      • NTR

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF:
  • FNAL:
  • GridPP:
  • IN2P3: Full day outage planned for December 8 (to be confirmed 1 week before)
  • JINR:
  • KISTI: NTR
  • KIT: NTR
  • NDGF:
  • NL-T1: NTR
  • NRC-KI:
  • OSG: Found 3 problem for the WLCG accounting, next report will be fixed. Still 1 problem to be solved for University of Florida.
  • PIC: Full day downtime on November 24 to perform dCache upgrade
  • RAL: NTR
  • TRIUMF: NTR

  • CERN batch and grid services:
    • myproxy.cern.ch's intervention has been rescheduled (Thu 19), please read the announcement placed on the ITSSB
    • LSF was down on Saturday due to a configuration problem after the recent migration from LSF 7 to LSF 9
  • CERN storage services:
  • Databases:
  • GGUS:
  • Grid Monitoring:
  • MW Officer:

AOB:

Thursday

Attendance:

  • local: Luca M. (SCOD+Storage), Maarten (ALICE), Fernando (Batch), Luca C. (DB)
  • remote: Christoph (CMS), Raja (LHCb), Michael (BNL), Rolf (IN2P3), Sang Un (KISTI), Thomas (KIT), Andrew (NL-T1), John (RAL), Kyle (OSG)

Experiments round table:

  • CMS reports (raw view) -
    • No major issues to report
    • Preparing for Heavy Ion run

  • ALICE -
    • also CERN and CNAF are ready for heavy ion data reconstruction
    • our thanks go to all experts involved!

  • LHCb reports (raw view) -
    • Data Processing:
      • Data processing of pp data at T0/T1/T2 sites.
      • Monte Carlo mostly at T0/T1/T2/T2D, user analysis at T0/1/2D sites
    • T0
      • BLAH error submitting to ce407 (GGUS:117694). Solved pretty fast - many thanks.
    • T1
      • CNAF : CVMFS error on various worker nodes (GGUS:117700)

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF:
  • FNAL:
  • GridPP:
  • IN2P3: NTR
  • JINR:
  • KISTI: NTR
  • KIT: NTR
  • NDGF:
  • NL-T1: NTR
  • NRC-KI:
  • OSG: NTR
  • PIC:
  • RAL: NTR
  • TRIUMF:

  • CERN batch and grid services:
    • On Monday 23/11/2015 the remaining ARC CE (ce501.cern.ch) will disappear from the site bdii.
      • The resource BDII direct from the CE will still be running.
      • The remaining VO who does not yet submit to HTCondorCE is encouraged to change their config to pull direct.
    • A CREAM issue on ce407 last night required a restart of tomcat.
  • CERN storage services: NTR
  • Databases: NTR
  • GGUS:
  • Grid Monitoring:
    • Final availability reports for October sent, and available at the SAM3 UI
  • MW Officer:

AOB:

  • Vidyo connection problem for Kyle (OSG)
Topic attachments
I Attachment History Action Size Date Who Comment
Unknown file formatpptx MB-Nov-15.pptx r1 manage 2857.5 K 2015-11-16 - 17:31 PabloSaiz  
Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2015-11-19 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback